1. 7x+6=3x+50
2. 6x+1=4x+15
3. 3x+2=-7+4x
4. 7x+10=26+5x

Answer by josephballard14(3)

7x+6=3x+50
-3x -3x from both sides
-4x+6=50
-6 -6 from both sides
-4x=44
divide both sides by -4
x=-11
6x+1=4x+15
-4x from both sides
2x+1=15
-1 from both sides
2x=14
divide 2 from both sides
x=7
3x+2=-7+4x
-4x from both sides
-1x+2=-7
-2 from both sides
-1x=-9
divide both sides by -1 and you get
x=9
7x+10=26+5x
-10 from both sides
7x=16+5x
-5x from both sides
2x=16
x=8